OpenStudy (lgbasallote):

What is the minimum number of coins you can use to come up with 67 cents out of quarters, dimes, nickels and pennies?

OpenStudy (anonymous):

67=25+25+10+5+1+1

OpenStudy (lgbasallote):

?

OpenStudy (zugzwang):

Six coins...?

OpenStudy (lgbasallote):

i can see that...but how is this solved?

OpenStudy (lgbasallote):

im pretty sure there's a better way than trial and error

OpenStudy (anonymous):

you have minimum numbers of coins, you have to put the possible largest "cents" first.

OpenStudy (anonymous):

Since 25>10>5>1, we have quarters first then 67-25=42>25, then we have another quarter 42-25=17 where 25>17>10, then we have a dime 17-10=7, where 10>7>5, then we have a nickel 7-5=2, so we have another 2 pennies. In total, 6 coins.

OpenStudy (lgbasallote):

ah that's a much better explanation than giving an answer and solution. won't you agree?

OpenStudy (anonymous):

I am lazy. :P

OpenStudy (zugzwang):

Wholeheartedly :) I'd just like to note that in a freak case, where they mint coins of denominations which are relatively prime, using up the "highest valued" coins may not always be best...

OpenStudy (zugzwang):

For instance, how to get 40 cents using 13, 15, and 1 cent coins

OpenStudy (anonymous):

I can show you an counter prove for "using up the "highest valued" coins may not always be best"

OpenStudy (zugzwang):

I thought you were lazy :) All the same, I'd be happy to see it

OpenStudy (anonymous):

We want to show that if we not using the "highest valued" coin first, we still can get minimum number of coins. Assume we have x cents with finite amount of coins. Suppose we have coins a>b>c and a>=b+c. If you do not use the "highest valued" a, then you have to use b and c instead of a which makes more amount of coins.